OpenStudy (anonymous):

Collect like terms. 31/9u+4/9v-17/9u+12/9v

OpenStudy (anonymous):

again lets group \[\frac{31}{9}u-\frac{17}{9}u=?\] looks bad, but it is not, because the denominators are the same

OpenStudy (anonymous):

and \[\frac{4}{9}v+\frac{12}{9}v=?\]

OpenStudy (anonymous):

\[\frac{31}{9}u-\frac{17}{9}u=\frac{31-17}{9}u=\frac{14}{9}u\] for the \(u\) part

OpenStudy (anonymous):

\[\frac{4}{9}v+\frac{12}{9}v=\frac{4+12}{9}v=\frac{16}{9}v\] for the \(v\)

OpenStudy (anonymous):

not too bad, right?

OpenStudy (anonymous):

so it would be \[\frac{ 14 }{ 9 }u+\frac{ 16 }{ 9 }v\]?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you, it really helped :)